OpenStudy (anonymous):

kenny spent one hour more than lucas studying for their chemistry test. andrea spent one hour less than twice the hours lucas spent studying for the test. if together the spent fourteen hours studying , how many hours did lucas study. translate into an equation.

OpenStudy (lgbasallote):

Let x = lucas's study time x+1 = kenny's study time x-1 = andrea's study time so x + x + 1 + x - 1= 14 so 3x = 14 x = 14/3

OpenStudy (anonymous):

Let x = lucas's study time x+1 = kenny's study time 2x-1 = andrea's study time so x + x + 1 + 2x - 1= 14 so 4x = 14 x = 14/4= 3 and 1/2 x+1=4 and 1/2 2x-1=2(3and1/2)-1 = 7-1 =6